1) CHRISTIAN IS 32 YEARS OLD AND HIS SON IS 5 YEARS OLD. HOW MANY YEARS FROM NOW WILL CHRISTIAN BE 10 TIMES HIS SON'S AGE.

Answer by srinivas(5)

There is no solution christian can never be 10 times his sons age.
32+x=(5+x)*10
32+x=50+10x
9x=-22
x=-22/9 which is absurd